# Large Language Models Share Representations of Latent Grammatical Concepts Across Typologically Diverse Languages

### v1

@@ Abstract
Jannik Brinkmann1,2   Chris Wendler3   Christian Bartelt1  Aaron Mueller2,4 1University of Mannheim  2Northeastern University  3EPFL  4Technion – IIT
@@ 1 Introduction
In the brains of human bilinguals, syntax processing may occur in similar regions for their first and second language, depending on factors like when the second language was learned (Cargnelutti et al., 2019), language proficiency (Połczyńska and Bookheimer, 2021), among many other factors (Sulpizio et al., 2020; Costa and Sebastián-Gallés, 2014). In multilingual language models (LMs; Shannon, 1948), how apt is the analogy of shared processing to human bilinguals? If we desire parameter-efficiency, we might want multilingual representations of concepts such as grammatical number, rather than many language-specific representations of the same concept.
@@
Past work has emphasized language-balanced pretraining corpora (e.g., Conneau et al., 2020; Xue et al., 2021; Mohammadshahi et al., 2022), such that an LM could be said to have many primary languages. However, many of the best-performing multilingual LMs are now primarily English models, trained on over 90% English text (e.g., Dubey et al., 2024). Why do these models perform so well in non-English languages? We hypothesize that these models learn generalizable abstractions that enable more efficient learning of new languages. In other words, being able to deeply characterize a smaller distribution could allow models to acquire more robust abstractions that may generalize more effectively to a wider distribution post-hoc; in contrast, more balanced corpora (wider distributions) could encourage the model to start by learning language-specific abstractions, which are potentially never merged into higher-level language-invariant concepts. For small models, on a synthetic language pair (English and a token-level duplication of it), Schäfer et al. (2024) show that imbalanced datasets lead to higher sample efficiency and higher overall performance in both languages in the low-data regime. [^1] For real languages, we expect that features will be shared across languages to a much greater extent.
@@
In large language models, multilinguality has both practical and theoretical scientific value. High performance on natural language processing tasks across many languages increases the impact and inclusivity of language technologies, and in a more compute-efficient manner than would be possible by training a large series of monolingual models. Moreover, multilingual models are hypothesized to be able to outperform monolingual models in low-resource languages: past work has emphasized the importance of cross-lingual transfer, where knowledge in one language is shared with another language whose corpora did not contain the information of interest. More abstractly, the extent to which abstract concepts are shared across languages addresses a key question: what kinds of multi-lingual processing mechanisms are acquirable from exposure to large-scale text distributions? It would be more parameter-efficient to learn generalizable abstractions that apply across languages, as opposed to redundantly learning the same grammatical feature for each language separately. Moreover, the existence of generalizable abstractions informs debates on memorization versus generalization: cross-linguistic generalization suggests a more sophisticated and broad application of particular concepts.
@@
Past work has largely investigated multilingual language models using behavioral/benchmarking analyses (Gupta and Srikumar, 2021; Raganato et al., 2020, inter alia) or neuron-level mechanistic analyses (Stanczak et al., 2022). Stanczak et al. observe that masked language models often shared neurons across languages for a particular concept, but causally verifying this is difficult, as counterfactual interventions to neurons often affect the processing of irrelevant concepts (see §2). To overcome this challenge, we make our units of causal analysis sparse autoencoder features, which have been shown to be more monosemantic, and therefore more human-interpretable (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024). This enables more precise causal interventions (Marks et al., 2024; Mueller et al., 2024), and allows us to qualitatively verify that features of interest are truly sensitive to the hypothesized feature, rather than a spurious but close-enough feature.
@@
In this work, we train a set of sparse autoencoders on the intermediate activations of Llama-3-8B Dubey et al. (2024) and Aya-3-8B (Aryabumi et al., 2024) and locate massively multilingual features for various morphosyntactic concepts. We design experiments to quantify the degree to which these concepts are shared across languages, and validate the their role in the models generations. Our results reveal that language models share morphosyntactic concept representations across typologically diverse languages, and that the internal lingua franca of large language models may not be English words per se, but rather concepts. [^2] Though we note that these concepts are likely to be biased toward English-like representations, as discussed in Wendler et al. (2024).
@@ 2 Background
The features underlying model computation are not guaranteed to be aligned to neuron bases; they may instead be represented in a distributed manner, such that there is a many-to-many relationship between neurons and concepts (Hinton et al., 1986; Smolensky, 1986). In practice, single neurons [^3] We use “neuron” to refer to a single dimension of any hidden representation vector in a model. are often polysemantic; in other words, they activate on a range of seemingly unrelated concepts (Bolukbasi et al., 2021; Elhage et al., 2022b). For example, Bricken et al. (2023) observe that a single neuron in a small LM responds to a mixture of academic citations, English dialogue, HTTP requests, and Korean text. To address this, Cunningham et al. (2024); Bricken et al. (2023) propose sparse autoencoders (SAEs) as a scalable technique for unsupervised discovery of interpretable feature directions in neural networks. Given an input activation ${\mathbf{x}}\in{\mathbb{R}}^{d_{\text{model}}}$, the autoencoder computes a decomposition
${\mathbf{x}}=\hat{{\mathbf{x}}}+{\bm{\epsilon}}(x)={\mathbf{b}}+\sum_{i}f_{i}({\mathbf{x}}){\mathbf{d}}_{i}+{\bm{\epsilon}}(x)$ (1)
into an approximate reconstruction $\hat{{\mathbf{x}}}$ as a linear combination of features ${\mathbf{d}}_{i}$. The features ${\mathbf{d}}_{i}\in{\mathbb{R}}^{d_{\text{model}}}$ are unit vectors, the feature activations $f_{i}({\mathbf{x}})\in{\mathbb{R}}$ are a sparse set of coefficients, ${\mathbf{b}}\in{\mathbb{R}}^{d_{\text{model}}}$ is a bias term, and ${\bm{\epsilon}}(x)\in{\mathbb{R}}^{d_{\text{model}}}$ is the approximation error.
Bricken et al. (2023) train SAEs by minimizing an L2 reconstruction error and an L1 regularization term to promote sparsity. However, the L1 regularization term introduces biases that can harm the accuracy of the reconstruction, as the loss can be decreased by trading-off reconstruction for lower L1. Therefore, we use the Gated SAE architecture (Rajamanoharan et al., 2024), which separates the functionality of selecting the features to use and estimating the activation magnitude of those directions. We provide additional details about Gated SAEs and their training in App. D.
A key idea underlying SAE-based interpretability is that one can reinterpret the language model’s (LM) internal computations in terms of the SAE features. By applying the decomposition in equation 1 to hidden states ${\mathbf{x}}$ in the LM and folding the SAEs into the forward pass, [^4] As described in detail in Marks et al. (2024). we can express model representations as a combination of SAE feature activations $f_{i}$ and reconstruction errors ${\bm{\epsilon}}$. This reframing enables attribution patching techniques (Syed et al., 2024) to identify which SAE features have causal influence on particular aspects of model behaviors.
@@
Causal interpretability methods aim to locate components in neural networks (such as attention heads or neurons) responsible for particular behaviors (Mueller et al., 2024). This requires constructing a distribution $\mathcal{D}$ over pairs of inputs $(x_{\text{clean}},x_{\text{patch}})$, where $x_{\text{clean}}$ is a prompt on which the behaviour occurs, and $x_{\text{patch}}$ is a reference prompt on which the behaviour does not occur. Then, we can quantify the importance of components using a target metric $\mathcal{L}:\mathbb{R}^{V}\rightarrow\mathbb{R}$, e.g., the difference in log-probabilities between a correct token continuation and minimally differing incorrect token continuation. In other words, we quantify the importance of a component on the model behaviour as the expected change in $\mathcal{L}$ when replacing the components’s activation on the clean prompt with its value on the patch prompt. This is known as the indirect effect (IE; Pearl, 2001).
An exact but compute-intensive method for causal analysis is activation patching (Meng et al., 2022), where we observe the change in model behaviour when performing a counterfactual intervention to a component. However, this requires a separate forward pass for each component. Therefore, we use attribution patching (AtP; Kramar et al., 2024) to estimate $\hat{\text{IE}}$, a linear approximation of the IE computed as a first-order Taylor expansion:
@@ 3.1 Experimental Setup
We select 23 languages from Universal Dependencies 2.1 (UD; Nivre et al., 2017), a multilingual treebank containing dependency-parsed sentences. This corresponds to the 23 languages that Aya-23 was trained on (see Appendix B). Each word in each sentence in UD is annotated with its part of speech and morphosyntactic features, as defined in the UniMorph schema (Kirov et al., 2018).
@@
We train Gated SAEs (Rajamanoharan et al., 2024) for both Llama-3-8B and Aya-23-8B. We collect activations over 250 million tokens of The Pile (Gao et al., 2020) extracted from the residual stream at the same layer at which we train the probing classifiers. To verify the quality of the SAEs, we measure the proportion of loss recovered when replacing the model activations with SAE reconstructions thereof. While the training dataset is primarily composed of English texts, we verify that it also reconstructs other language effectively using the mC4 dataset (Xue et al., 2021; see App. D).
@@ 3.2 Results
Representations of morphosynthactic concepts are shared across languages.
First, we measure the extent of overlap across languages for the top features encoding a morphosyntactic category. For this, we perform attribution patching (§2), where the target metric is the logit of the probing classifiers we trained in §2. This identifies the set of causally relevant features that are most influential on the probe’s logits. We observe that a small fraction of features dominate the causal effect, with a long tail of features that have a small effect. Therefore, for each language and concept, we select the top-32 most informative features per probe and compute the overlap across languages within a concept. We present the results in Figure 2.
We observe significant overlap of up to 50 % in the features identified for the same grammatical concepts across typologically diverse languages. For example, feature 22860 is among the most influential features across all 15 languages that inflect for grammatical gender. The overlap is greater for morphosyntactic concepts present in a larger number of languages, such as grammatical number. We also find that the extent of overlap is largely consistent across Llama-3-8B and Aya-23-8B, with only a small number of exceptions (e.g. negative polarity). Thus, despite the lack of one-to-one mapping of grammatical concepts across languages, there exist highly multilingual representations of abstract morphosyntactic concepts.
@@
Figure 3: Examples of the activation patterns of selected features that correspond to cross-lingual representations of grammatical concepts. For example, we locate features that indicate the presence of plural nouns across languages or features that indicate past tense.
Figure 4: Performance of the probing classifiers before and after ablating multilingual features. Specifically, we test ablating (a) all multilingual features, and (b) the upper quartile of the most multilingual features. We find that the classifiers crucialy rely on multilingual features to predict the presence of morphosyntactic features.
@@
We also observe that while many features are shared across at least two languages, there is a small set of features shared across a large number of languages (App. E Therefore, we investigate ablating only massively multilingual features. Specifically, we order features by the number of languages for which they are causally relevant, and the average strength of their causal effect. Then, we select the upper quartile of features in this ranking—i.e., the most multilingual features—and only ablate those. For Llama 3, the performance after ablating only massively multilingual features drops to 67% on average. Notably, when ablating all multilingual features, the performance drops to 61%—a difference of only 6% percent, despite ablating four times as many features. This suggests that massively multilingual features explain most of the probes’ behavior.
@@ 4 Demonstrating Functional Selectivity: A Case Study in Machine Translation
Another benefit of sparse features is the level of fine-grained control they provide over language model behaviors (Templeton et al., 2024). Indeed, feature-based interventions tend to be more effective then neuron-based methods at precisely modifying how language models generalize without destroying task performance (Marks et al., 2024). We leverage this property in a machine translation setting to (i) demonstrate the functional selectivity of the discovered morphosyntactic features (i.e., to show that intervening with them has few side effects on other concepts), and (ii) provide additional evidence that their causal role is highly general (i.e., that it is effective in this MT setting and not just on the probes). To evaluate (i) and (ii), we use feature steering (Templeton et al., 2024), where we clamp features of interest to counterfactual activations. These activations can be either significantly higher or lower than the values observed in practice. This approach is similar to prior work on understanding the functional role of individual neurons in image models (Bau et al., 2019).
As translation is a difficult task that does not always have a single clear answer, we restrict to a smaller subsample of concepts and languages such that we can manually verify the baseline quality of translations, and then the efficacy of the interventions. We restrict our analysis to tense (Past vs. Present), gender (Masc. vs. Fem.), and number (Sing. vs. Plural), as these concepts are shared across most languages we consider, and narrow to three languages: English, German, and French. We provide details on the prompt format and translation performance in Appendix F.

### v2

@@ Abstract
Jannik Brinkmann1,2     Chris Wendler3     Christian Bartelt2    Aaron Mueller1,4 1Northeastern University  2TU Clausthal  3EPFL  4Technion – IIT
@@ 1 Introduction
In the brains of human bilinguals, syntax processing may occur in similar regions for their first and second language, depending on factors like when the second language was learned (Cargnelutti et al., 2019), language proficiency (Połczyńska and Bookheimer, 2021), among many other factors (Sulpizio et al., 2020; Costa and Sebastián-Gallés, 2014). In multilingual language models (LMs; Shannon, 1948), how apt is the analogy of shared processing to human bilinguals? If we desire parameter-efficiency, we might want multilingual representations of concepts such as grammatical number, rather than many language-specific representations of the same concept. We release code, data, and autoencoders at github.com/jannik-brinkmann/multilingual-features.
@@
Past work has emphasized language-balanced pretraining corpora (e.g., Conneau et al., 2020; Xue et al., 2021; Mohammadshahi et al., 2022), such that an LM could be said to have many primary languages. However, many of the best-performing multilingual LMs are now primarily English models, trained on over 90% English text (e.g., Dubey et al., 2024). Why do these models perform so well in non-English languages? We hypothesize that these models learn generalizable abstractions that enable more efficient learning of new languages. In other words, being able to deeply characterize a smaller distribution could allow models to acquire more robust abstractions that may generalize more effectively to a wider distribution post-hoc; in contrast, more balanced corpora (wider distributions) could encourage the model to start by learning language-specific abstractions, which are potentially never merged into higher-level language-invariant concepts. For small models, on a synthetic language pair (English and a token-level duplication of it), Schäfer et al. show that imbalanced datasets lead to higher sample efficiency and higher overall performance in both languages in the low-data regime. [^1] For real languages, we expect that features will be shared across languages to a much greater extent.
@@
In large language models, multilinguality has both practical and theoretical scientific value. High performance on natural language processing tasks across many languages increases the impact and inclusivity of language technologies, and in a more compute-efficient manner than would be possible by training a large series of monolingual models. Moreover, multilingual models are hypothesized to be able to outperform monolingual models in low-resource languages: past work has emphasized the importance of cross-lingual transfer, where knowledge in one language is shared with another language whose corpora did not contain the information of interest (Libovický et al., 2020; Chang et al., 2022; Hua et al., 2024). More abstractly, the extent to which abstract concepts are shared across languages addresses a key question: what kinds of multi-lingual processing mechanisms are acquirable from exposure to large-scale text distributions? It would be more parameter-efficient to learn generalizable abstractions that apply across languages, as opposed to redundantly learning the same grammatical feature for each language separately. Moreover, the existence of generalizable abstractions informs debates on memorization versus generalization: cross-linguistic generalization suggests a more sophisticated and broad application of particular concepts.
@@
Past work has largely investigated multilingual language models using behavioral/benchmarking analyses (Gupta and Srikumar, 2021; Raganato et al., 2020, inter alia) or neuron-level mechanistic analyses (Stanczak et al., 2022; de Varda and Marelli, 2023). Stanczak et al. observe that masked language models often share neurons across languages for a particular concept, but causally verifying this is difficult, as counterfactual interventions to neurons often affect the processing of irrelevant concepts (see §2). To overcome this challenge, we make our units of causal analysis sparse autoencoder features, which have been shown to be more monosemantic, and therefore more human-interpretable (Cunningham et al., 2024; Bricken et al., 2023; Rajamanoharan et al., 2024). This enables more precise causal interventions (Marks et al., 2025; Mueller et al., 2024), and allows us to qualitatively verify that features of interest are truly sensitive to the hypothesized feature, rather than a spurious but close-enough feature.
@@
In this work, we train a set of sparse autoencoders on the intermediate activations of Llama-3-8B Dubey et al. (2024) and Aya-3-8B (Aryabumi et al., 2024) and locate (massively) multilingual features for various morphosyntactic concepts. We design experiments to quantify the degree to which these concepts are shared across languages, and validate their role in the models generations. Our results reveal that language models share morphosyntactic concept representations across typologically diverse languages, and that the internal lingua franca of large language models may not be English words per se, but rather concepts. [^2] Though we note that these concepts are likely to be biased toward English-like representations, as discussed in Wendler et al. (2024).
@@ 2 Background
The features underlying model computation are not guaranteed to be aligned to neuron bases; they may instead be represented in a distributed manner, such that there is a many-to-many relationship between neurons and concepts (Hinton et al., 1986; Smolensky, 1986). In practice, single neurons [^3] We use “neuron” to refer to a single dimension of any hidden representation vector in a model. are often polysemantic; in other words, they activate on a range of seemingly unrelated concepts (Bolukbasi et al., 2021; Elhage et al., 2022b). For example, Bricken et al. (2023) observe that a single neuron in a small LM responds to a mixture of academic citations, English dialogue, HTTP requests, and Korean text. To address this, Cunningham et al. (2024); Bricken et al. (2023) propose sparse autoencoders (SAEs) as a scalable technique for unsupervised discovery of interpretable feature directions in neural networks. Given an activation vector ${\mathbf{x}}\in{\mathbb{R}}^{d_{\text{model}}}$, the autoencoder computes a decomposition
${\mathbf{x}}=\hat{{\mathbf{x}}}+{\bm{\epsilon}}(x)={\mathbf{b}}+\sum_{i}f_{i}({\mathbf{x}}){\mathbf{v}}_{i}+{\bm{\epsilon}}(x)$ (1)
into an approximate reconstruction $\hat{{\mathbf{x}}}$ as a linear combination of features ${\mathbf{v}}_{i}$. The features ${\mathbf{v}}_{i}\in{\mathbb{R}}^{d_{\text{model}}}$ are unit vectors, the feature activations $f_{i}({\mathbf{x}})\in{\mathbb{R}}$ are a sparse set of coefficients, ${\mathbf{b}}\in{\mathbb{R}}^{d_{\text{model}}}$ is a bias term, and ${\bm{\epsilon}}(x)\in{\mathbb{R}}^{d_{\text{model}}}$ is the approximation error.
Bricken et al. (2023) train SAEs by minimizing an L2 reconstruction error and an L1 regularization term to promote sparsity. However, the L1 regularization term introduces biases that can harm the accuracy of the reconstruction. Therefore, we use the Gated SAE architecture (Rajamanoharan et al., 2024), which separates the functionality of selecting the features to use and estimating the activation magnitude. We provide additional details about Gated SAEs and their training in App. D.
A key idea underlying SAE-based interpretability is that one can reinterpret the language model’s (LM) internal computations in terms of the SAE features. By applying the decomposition in Eq. 1 to hidden states ${\mathbf{x}}$ in the LM and folding the SAEs into the forward pass, [^4] As described in detail in Marks et al. (2025). we can express model representations as a combination of SAE feature activations $f_{i}$ and reconstruction errors ${\bm{\epsilon}}$. This reframing enables gradient-based attribution patching techniques (Syed et al., 2024) to identify which SAE features have causal influence on particular aspects of model behavior.
@@
Causal interpretability methods aim to locate components in neural networks (such as attention heads or neurons) responsible for particular behaviors (Mueller et al., 2024). This requires constructing a distribution $\mathcal{D}$ over pairs of inputs $(x_{\text{clean}},x_{\text{patch}})$, where $x_{\text{clean}}$ is a prompt on which the behaviour occurs, and $x_{\text{patch}}$ is a reference prompt on which the behaviour does not occur. Then, we can quantify the importance of components using a target metric $m:\mathbb{R}^{V}\rightarrow\mathbb{R}$, e.g., the difference in log-probabilities between a correct token continuation and minimally differing incorrect token continuation. In other words, we quantify the importance of a component on the model behaviour as the expected change in $\mathcal{L}$ when replacing the components’s activation on the clean prompt with its value on the patch prompt. This is known as the indirect effect (IE; Pearl, 2001).
An exact but compute-intensive method for causal analysis is activation patching (Vig et al., 2020), where we observe the change in model behaviour when performing a counterfactual intervention to a component. However, this requires a separate forward pass for each component. Therefore, we use attribution patching (AtP; Kramar et al., 2024) to estimate $\hat{\text{IE}}$, a linear approximation of the IE computed as a first-order Taylor expansion:
@@ 3.1 Experimental Setup
We select 23 languages from Universal Dependencies 2.1 (UD; Nivre et al., 2017), a multilingual treebank containing dependency-parsed sentences. This corresponds to the 23 languages that Aya-23 was trained on (see App. B). Each word in each sentence in UD is annotated with its part of speech and morphosyntactic features, as defined in the UniMorph schema (Kirov et al., 2018).
@@
We train Gated SAEs (Rajamanoharan et al., 2024) for both Llama-3-8B and Aya-23-8B. We collect activations over 250 million tokens of The Pile (Gao et al., 2020) extracted from the residual stream at the same layer at which we train the probing classifiers. To verify the quality of the SAEs, we measure the proportion of loss recovered when replacing the model activations with SAE reconstructions thereof. While the training dataset is primarily composed of English texts, we verify that it also reconstructs other languages effectively using the mC4 dataset (Xue et al., 2021; see App. D).
@@ 3.2 Results
Representations of morphosyntactic concepts are shared across languages.
First, we measure the extent of overlap across languages for the top features encoding a morphosyntactic category. For this, we perform attribution patching (§2), where the target metric is the logit of the probing classifiers we trained in §3.1. This identifies the set of causally relevant features that are most influential on the probe’s predictions. We observe that a small fraction of features dominate the causal effect, with a long tail of features that have a small effect. Therefore, for each language and concept, we select the top-32 most informative features per probe and compute the overlap across languages within a concept. We present the results in Figure 2.
We observe significant overlap of up to 50% in the features identified for the same grammatical concepts across typologically diverse languages. For example, feature 22860 is among the most influential features across all 15 languages that inflect for grammatical gender. The overlap is greater for morphosyntactic concepts present in a larger number of languages, such as grammatical number. We also find that the extent of overlap is largely consistent across Llama-3-8B and Aya-23-8B, with only a small number of exceptions (e.g. negative polarity). Thus, despite the lack of a one-to-one mapping of grammatical concepts across languages, there exist highly multilingual representations of abstract morphosyntactic concepts.
@@
Figure 3: Examples of the activation patterns of selected features in Llama-3-8B that correspond to cross-lingual representations of grammatical concepts. For example, we locate features that indicate the presence of plural nouns or features that indicate past tense across languages.
Figure 4: Performance of the probing classifiers before and after ablating features. Specifically, we test ablating (a) all monolingual features, (b) all multilingual features, and (c) the upper quartile of the most multilingual features. We find that the classifiers crucialy rely on massively multilingual features to predict the presence of morphosyntactic features.
@@
We also observe that while many features are shared across at least two languages, there is a small set of features shared across a large number of languages (App. E). Therefore, we investigate ablating only massively multilingual features. Specifically, we order features by the number of languages for which they are causally relevant, and the average strength of their causal effect. Then, we select the upper quartile of features in this ranking—i.e., the most multilingual features—and only ablate those. For Llama 3, the performance after ablating only massively multilingual features drops to 64% on average. Notably, when ablating all multilingual features, the performance drops to 61%—a difference of only 3%, despite ablating four times as many features. This suggests that the most massively multilingual features explain most of the probes’ behavior.
@@ 4 Demonstrating Functional Selectivity: A Case Study in Machine Translation
Another benefit of sparse features is the level of fine-grained control they provide over language model behaviors (Templeton et al., 2024). Indeed, feature-based interventions tend to be more effective then neuron-based methods at precisely modifying how language models generalize without destroying task performance (Marks et al., 2025). We leverage this property in a machine translation setting to (i) demonstrate the functional selectivity of the discovered morphosyntactic features (i.e., to show that intervening with them has few side effects on other concepts), and (ii) provide additional evidence that their causal role is highly general (i.e., that it is effective in this MT setting and not just on the probes). To evaluate (i) and (ii), we use feature steering (Templeton et al., 2024), where we clamp features of interest to counterfactual activations. These activations can be either significantly higher or lower than the values observed in practice. This approach is similar to prior work on understanding the functional role of individual neurons in image models (Bau et al., 2019).
Figure 5: Efficacy in flipping the model behavior on our dataset when translating while intervening on a single multilingual feature per concept. For each concept, we translate a sentence containing some concept (e.g., present tense) from some source language to another language while intervening on a feature, and measure the number of times the model generates a translation containing the counterfactual concept value (e.g. past tense). In each setting, we intervene on a single feature and measure the success rate over 64 examples. Results aggregated across translation directions; see App. F.3 for separate results per direction.
As translation is a difficult task that does not always have a single clear answer, we restrict to a smaller subsample of concepts and languages such that we can manually verify the baseline quality of translations, and then the efficacy of the interventions. We restrict our analysis to tense (Past vs. Present), gender (Masc. vs. Fem.), and number (Sing. vs. Plural), as these concepts are shared across most languages we consider. We analyze English, French, and German; we additionally include Turkish, but restrict our focus to number and tense here, as Turkish does not inflect for grammatical gender. We provide details on the prompt format and translation performance in Appendix F.